### v2

Abstract
The Betti tables of a multigraded module encode the grades at which there is an algebraic change in the module. Multigraded modules show up in many areas of pure and applied mathematics, and in particular in topological data analysis, where they are known as persistence modules, and where their Betti tables describe the places at which the homology of filtered simplicial complexes changes. Although Betti tables of singly and bigraded modules are already being used in applications of topological data analysis, their computation in the bigraded case (which relies on an algorithm that is cubic in the size of the filtered simplicial complex) is a bottleneck when working with large datasets. We show that, in the special case of 00-dimensional homology (relevant for clustering and graph classification) Betti tables of bigraded modules can be computed in log-linear time. We also consider the problem of computing minimal presentations, and show that minimal presentations of 00-dimensional persistent homology can be computed in quadratic time, regardless of the grading poset.
keywords:
Multiparameter persistence, Zero-dimensional homology, Minimal presentation, Betti table
1 Introduction
Betti tables and persistence
Betti tables are a classical descriptor of a multigraded modules [19, 30, 34], which encode the grades of the generators in a minimal projective resolution of the module (see, e.g., Fig. 1 and Example 2.15). Informally, one can interpret the Betti tables of a graded module as recording the grades at which there is an algebraic change in the module. Graded modules have applications in a wide variety of areas of pure and applied mathematics, including topological data analysis, and more specifically, persistence theory [33, 6], where they are known as persistence modules, and where they are used to describe the varying topology of simplicial complexes and other spaces as they are filtered by one or more real parameters.
Informally, one-parameter persistence modules correspond to $\mathbb{Z}$-graded $\mathbb{k}[x]$-modules, and can thus be classified up to isomorphism effectively. Multiparameter persistence modules [11, 6] correspond to $\mathbb{Z}^{n}$-graded $\mathbb{k}[x_{1},\dots,x_{n}]$-modules, and thus do not admit any reasonable classification up to isomorphism (formally, one is dealing with categories of wild representation type [31]; see [11, 2, 4] for manifestations of this phenomenon in persistence theory). For this reason, much of the research in multiparameter persistence is devoted to the study of incomplete descriptors of multiparameter persistence modules. Betti tables (also known as multigraded Betti numbers) provide one of the simplest such descriptors, and various properties of this descriptor from the point of view of persistence theory are well understood, including their effective computation [28, 26, 20, 3], their relationship to discrete Morse theory [23, 1, 22], their optimal transport Lipschitz-continuity with respect to perturbations [32], and their usage in supervised learning [29, 39].
Two-parameter persistent homology
[…]
Figure 1: Left. A bifiltered graph $(G,f)$ with vertex set $\{u,v,w,x_{1},x_{2},x_{3}\}$ and edge set $\{e_{1},e_{2},e_{3},h_{1},h_{2},d_{1},d_{2},d_{3}\}$. Right. The bifiltered graph schematically mapped to $\mathbb{R}^{2}$, together with the Betti tables $\beta_{0}(H_{0}(G,f))$ (circles), $\beta_{1}(H_{0}(G,f))$ (crosses), $\beta_{2}(H_{0}(G,f))$ (stars), and $\beta_{0}(H_{1}(G,f))$ (squares).
The current standard algorithm for computing the Betti tables of $H_{i}(K,f;\mathbb{k})$ in the two-parameter case is the Lesnick–Wright algorithm [28], which runs in time $O(|K|^{3})$. Because of results such as those of [18], one does not expect to find algorithms with better worst-case time complexity than matrix-multiplication time, at least when $i\in\mathbb{N}$ is arbitrary. Current options to speed up practical computations include sparsifying the filtered complex before computing homology [42], as well as computational shortcuts that are known to significantly reduce computational time in practice [26, 3]. Nevertheless, computational cost is still a main bottleneck in real-world applications of persistence, limiting the size of the datasets on which it can be applied.
Zero-dimensional persistent homology
In many applications, persistent homology in dimension zero is all that is required, as it encodes information about the changes in connectivity of filtered simplicial complexes, making it useful for clustering [10, 9, 15, 8, 35, 38] and graph classification [41, 13, 25, 29, 24].
But, if one is only interested in 00-dimensional homology $H_{0}(K,f;\mathbb{k})$, algorithms relying on linear algebra are usually far from the most efficient ones: For example, in the one-parameter case, the Betti tables of the 00-dimensional homology of an $\mathbb{R}$-filtered graph $(G,f)$ can be computed in $O(|G|\,\log|G|)$ time by first sorting the simplices of $G$ by their $f$-value, and then doing an ordered pass using a union-find data structure; in the language of barcodes, the Betti tables simply record the endpoints of the barcode, and the barcode can be computed using the elder rule (see [17, pp. 188] or [35, Algorithm 1]).
Contributions
The paper is concerned with the following question: What is the complexity of computing the Betti tables of graphs filtered by posets other than $\mathbb{R}$?
We introduce algorithms for the computation of a minimal set of generators and of a minimal presentation of 00-dimensional persistent homology indexed by an arbitrary poset.
Theorem A.
Let $\mathscr{P}$ be any poset, and let $(G,f)$ be a finite $\mathscr{P}$-filtered graph. Algorithm 2 computes the 00th Betti table of $H_{0}(G,f;\mathbb{k})$ in $O\big{(}\,|G|\,\big{)}$ time. Algorithm 3 computes a minimal presentation (and hence the 00th and $1$st Betti tables) of $H_{0}(G,f;\mathbb{k})$ in $O(\,|G|^{2}\,)$ time.
We also introduce a more efficient algorithm specialized to the two-parameter case, which computes all Betti tables, that is, 00th, $1$st, and $2$nd, by Hilbert’s syzygy theorem.
Theorem B.
Let $(G,f)$ be a finite $\mathbb{R}^{2}$-filtered graph. Algorithm 4 computes minimal presentations and the Betti tables of $H_{0}(G,f;\mathbb{k})$ and $H_{1}(G,f;\mathbb{k})$ in $O(\,|G|\,\log|G|\,)$ time.
Of note is the fact that Betti tables of 00-dimensional two-parameter persistent homology can be computed in log-linear time, as in the one-parameter case.
As another main contribution, we establish a connection between minimal presentations and connectivity properties of filtered graphs (Theorems 3.8 and 3.10), which allows us to abstract away the algebraic problem, and focus on a simpler combinatorial problem. The correctness of our algorithms is based on this. Another interesting consequence of these results is that, for arbitrary posets, the 00th and $1$st Betti tables of 00-dimensional persistent homology are independent of the field of coefficients, while for the poset $\mathbb{R}^{2}$, all Betti tables of a filtered graph are independent of the field of coefficients.
Summary of approach and structure of the paper
The main body of the paper has three sections: one on background (Section 2), one on theoretical results (Section 3), and one on algorithms (Section 4). Appendix A contains most proofs.
In order to describe and prove the correctness of our algorithms, we introduce, in the theory section, the notion of a minimal filtered graph (Definition 3.2). Informally, a minimal filtered graph is one whose vertices and edges induce a minimal presentation of its 00-dimensional persistent homology (see Theorem 3.8). A graph is not minimal if it has some edge that can be either contracted or deleted without changing its 00-dimensional persistent homology (for example, the graph of Fig. 1 has both contractible and deletable edges, and is thus not minimal; see Examples 2.15 and 3.3). The idea is that, by contracting and deleting edges that do not change 00-dimensional persistent homology, one inevitably ends up with a minimal filtered graph, from which a minimal presentation can be easily extracted. Our main theoretical contributions (Theorems 3.8 and 3.10) make this idea precise by giving an explicit minimal presentation of $H_{0}$ of any minimal filtered graph, and an explicit minimal resolution of $H_{0}$ of any minimal $\mathbb{R}^{2}$-filtered graph. Our main algorithmic contributions are efficient algorithms for contracting and deleting edges as necessary. Algorithm 4 makes use of a dynamic tree data structure [40], which is the main ingredient that allows us to compute the Betti tables of bifilitered graphs in log-linear time; we give more details in Section 4.
Remark about multi-critical filtrations
The filtrations considered in this paper are $1$-critical, meaning that each simplex (vertex or edge, in the case of graphs) appears at exactly one grade. In Appendix B, we describe a simple preprocessing step to turn a finite multi-critical filtration by a lattice into a $1$-critical filtration with the same 00-dimensional homology. In the two-parameter case, this construction does not change the input complexity, but for other lattices it may increase the input size. Note that the construction does not (necessarily) preserve $1$-dimensional homology.
Related work
To the best of our knowledge, the only subcubic algorithm related to Algorithm 4 is that of [8], which, in particular, can be used to compute the Betti tables of $H_{0}$ of a function-Rips complex of a finite metric space $X$ in time $O(|X|^{2}\,\log|X|)$. When applied to a function-Rips complex, our Algorithm 4 has the same time complexity; however, our Algorithm 4 applies to arbitrary bifiltered graphs, while function-Rips complexes are very special (and do not include arbitrary filtered graphs). The inner workings of Algorithm 4 are also different from those of [8]: While we rely on dynamic trees, their algorithm relies on a dynamic minimum spanning tree, and, notably, on the fact that, in a function-Rips bifiltration, vertices are filtered exclusively by one of the two filtering functions (so that vertices are linearly ordered in the bifiltration).
The computation of minimal presentations of $\mathbb{R}^{n}$-filtered complexes is studied in [5]. Since they consider homology in all dimensions, their complexity is significantly worse than the quadratic complexity of Algorithm 3, which only applies to 00-dimensional homology.
Part of the contributions in this paper could be rephrased using the language of discrete Morse theory for multiparameter filtrations [1, 37, 7], specifically, Algorithms 1 and 2 are essentially computing acyclic matchings which respect the filtration. However, this point of view requires extra background, and, to the best of our understanding, cannot be used to describe the more interesting Algorithm 4.
2 Background
As is common in persistence theory, we assume familiarity with very basic notions of category theory, specifically, that of a category and of a functor. We let $\mathbb{k}$ denote a field, $\mathbf{Vec}_{\mathbb{k}}$ denote the category of $\mathbb{k}$-vector spaces, and $\mathbf{Set}$ denote the category of sets. When the field $\mathbb{k}$ plays no role, we may denote $\mathbf{Vec}_{\mathbb{k}}$ simply by $\mathbf{Vec}$. Proofs can be found in Section A.1.
Graphs and filtered graphs
A graph $G=(V,E,\partial)$ consists of finite sets $V$ and $E$, and a function $\partial:E\xrightarrow{\;\;\;}V\times V$. We refer to the elements of $V$ as vertices, typically denoted $v,w,x,y\in V$, and to the elements of $E$ as edges, typically denoted $e,d,h\in E$. If $\partial e=(v,w)$, we write $e_{0}=v$ and $e_{1}=w$. The size of a graph $G$ is $|G|=|V|+|E|$.
A subgraph of a graph $G=(V,E,\partial)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime},\partial^{\prime})$, where $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$, and such that $\partial^{\prime}=\partial|_{E^{\prime}}$ takes values in $V^{\prime}\times V^{\prime}\subseteq V\times V$. If $G^{\prime}$ is a subgraph of $G$, we write $G^{\prime}\subseteq G$.
If $E^{\prime}\subseteq E$ is a set of edges of $G$, we let $G\setminus E^{\prime}$ be the subgraph of $G$ with the same vertices and $E\setminus E^{\prime}$ as set of edges.
Let $\mathscr{P}$ be a poset. A $\mathscr{P}$-filtered graph $(G,f^{V},f^{E})$ consists of a graph $G$ and functions $f^{V}:V\xrightarrow{\;\;\;}\mathscr{P}$ and $f^{E}:E\xrightarrow{\;\;\;}\mathscr{P}$ such that $f^{V}(e_{0})\leq f^{E}(e)$ and $f^{V}(e_{1})\leq f^{E}(e)$ for all $e\in E$. When there is no risk of confusion, we refer to $\mathscr{P}$-filtered graphs simply as filtered graph, and denote both $f^{V}$ and $f^{E}$ by $f$ and the filtered graph $(G,f^{V},f^{E})$ by $(G,f)$.
If $(G,f)$ is a $\mathscr{P}$-filtered graph and $r\in\mathscr{P}$, we let $(G,f)_{r}$ be the subgraph of $G$ with vertices $\{v\in V:f(v)\leq r\}$ and edges $\{e\in E:f(e)\leq r\}$.
Persistence modules and persistent sets
Let $\mathscr{P}$ be a poset. A $\mathscr{P}$-persistence module is a functor $\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$, where $\mathscr{P}$ is the category associated with $\mathscr{P}$. Explicitly, a $\mathscr{P}$-persistence module $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ consists of the following:
  • for each $r\in\mathscr{P}$, a vector space $M(r)$;
  • for each pair $r\leq s\in\mathscr{P}$, a linear morphism $\varphi^{M}_{r,s}:M(r)\xrightarrow{\;\;\;}M(s)$; such that
  • for all $r\in\mathscr{P}$, the linear morphism $\varphi^{M}_{r,r}:M(r)\xrightarrow{\;\;\;}M(r)$ is the identity;
  • for all $r\leq s\leq t\in\mathscr{P}$, we have $\varphi^{M}_{s,t}\circ\varphi^{M}_{r,s}=\varphi^{M}_{r,t}:M(r)\xrightarrow{\;\;\;}M(t)$.
When there is no risk of confusion, we may refer to a $\mathscr{P}$-persistence module as a persistence module. An $n$-parameter persistence module ($n\geq 1\in\mathbb{N}$) is an $\mathbb{R}^{n}$-persistence module.
A morphism $g:M\xrightarrow{\;\;\;}N$ between persistence modules is a natural transformation between functors, that is, a family of linear maps $\{g_{r}:M(r)\xrightarrow{\;\;\;}N(r)\}_{r\in\mathscr{P}}$ with the property that $\varphi^{N}_{r,s}\circ g_{r}=g_{s}\circ\varphi^{M}_{r,s}:M(r)\xrightarrow{\;\;\;}N(s)$, for all $r\leq s\in\mathscr{P}$. Such a morphism is an isomorphism if $g_{r}:M(r)\xrightarrow{\;\;\;}N(r)$ is an isomorphism of vector spaces for all $r\in\mathscr{P}$.
If $M,N:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ are persistence modules, their direct sum, denoted $M\oplus N:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$, is the persistence module with $(M\oplus N)(r)\coloneqq M(r)\oplus N(r)$ and with $\varphi^{M\oplus N}_{r,s}\coloneqq\varphi^{M}_{r,s}\oplus\varphi^{N}_{r,s}:M(r)\oplus N(r)\xrightarrow{\;\;\;}M(s)\oplus N(s)$, for all $r\leq s\in\mathscr{P}$.
Similarly, a $\mathscr{P}$-persistent set is a functor $\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Set}$. The concepts of $n$-parameter persistent set, and of morphism and isomorphism between persistent sets are defined analogously.
Persistent homology and connected components of filtered graphs
If $S\in\mathbf{Set}$, we let $\langle S\rangle_{\mathbb{k}}\in\mathbf{Vec}$ denote the free vector space generated by $S$; this defines a functor $\langle-\rangle_{\mathbb{k}}:\mathbf{Set}\xrightarrow{\;\;\;}\mathbf{Vec}_{\mathbb{k}}$. Let $G=(V,E,\partial)$ be a graph. Consider the $\mathbb{k}$-linear map
[…]
$e$ $\xmapsto{\;\;\;\;\;\;}e_{1}-e_{0}$
The 00-dimensional homology of $G$, denoted $H_{0}(G;\mathbb{k})$, is the $\mathbb{k}$-vector space $\operatorname{coker}(d)$, and the $1$-dimensional homology of $G$, denoted $H_{1}(G;\mathbb{k})$, is the $\mathbb{k}$-vector space $\operatorname{ker}(d)$. In particular, every vertex $v\in V$ gives an element $[v]\in H_{0}(G;\mathbb{k})$. When there is no risk of confusion, we omit the field $\mathbb{k}$ and write $H_{i}(G)$ instead of $H_{i}(G;\mathbb{k})$.
Homology is functorial, in the following sense. If $G^{\prime}=(V^{\prime},E^{\prime},\partial^{\prime})$ is a subgraph of $G$, we have a commutative square
[…]
induced by the inclusions $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. This induces linear maps $H_{0}(G^{\prime})\xrightarrow{\;\;\;}H_{0}(G)$ and $H_{1}(G^{\prime})\xrightarrow{\;\;\;}H_{1}(G)$. Moreover, the morphism $H_{\bullet}(G^{\prime\prime})\xrightarrow{\;\;\;}H_{\bullet}(G)$ induced by a subgraph $G^{\prime\prime}\subseteq G^{\prime}\subseteq G$ is equal to the composite $H_{\bullet}(G^{\prime\prime})\xrightarrow{\;\;\;}H_{\bullet}(G^{\prime})\xrightarrow{\;\;\;}H_{\bullet}(G)$.
If $(G,f)$ is a $\mathscr{P}$-filtered graph, and $i\in\{0,1\}$, we get a $\mathscr{P}$-persistence module $H_{i}(G,f):\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$, with $H_{i}(G,f)(r)=H_{i}((G,f)_{r})$, and with structure morphism $H_{i}(G,f)(r)\xrightarrow{\;\;\;}H_{i}(G,f)(s)$ for $r\leq s\in\mathscr{P}$ induced by the inclusion of graphs $(G,f)_{r}\subseteq(G,f)_{s}$.
The set of connected components of $G$, denoted $\pi_{0}(G)$, is the quotient of $V$ by the equivalence relation $\sim$ where $v\sim w\in V$ if and only if there exists a path in $G$ between $v$ and $w$. If $v\in V$, we let $[v]\in\pi_{0}(G)$ denote its connected component, so that $[v]=[w]\in\pi_{0}(G)$ if and only if $v$ and $w$ belong to the same connected component.
The set of connected components is also functorial with respect to inclusions $G^{\prime}\subseteq G$, since $[v]=[w]\in\pi_{0}(G^{\prime})$ implies $[v]=[w]\in\pi_{0}(G)$. In particular, if $(G,f)$ is a $\mathscr{P}$-filtered graph, we get a $\mathscr{P}$-persistent set $\pi_{0}(G,f):\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Set}$, with $\pi_{0}(G,f)(r)=\pi_{0}((G,f)_{r})$, and with the structure morphism $\pi_{0}(G,f)(r)\xrightarrow{\;\;\;}\pi_{0}(G,f)(s)$ for $r\leq s\in\mathscr{P}$ induced by the inclusion of graphs $(G,f)_{r}\subseteq(G,f)_{s}$.
The following is straightforward to check.
Lemma 2.1.
If $G$ is a graph, then the map $\langle\pi_{0}(G)\rangle_{\mathbb{k}}\xrightarrow{\;\;\;}H_{0}(G)$ sending a basis element $[v]\in\pi_{0}(G)$ to $[v]\in H_{0}(G)$ is well-defined and an isomorphism of vector spaces. In particular, if $(G,f)$ is a $\mathscr{P}$-filtered graph, composing the persistent set $\pi_{0}(G,f):\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Set}$ with the free vector space functor $\langle-\rangle_{\mathbb{k}}:\mathbf{Set}\xrightarrow{\;\;\;}\mathbf{Vec}$ yields a persistence module isomorphic to $H_{0}(G,f):\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$. ∎
Projective persistence modules
Given $r\in\mathscr{P}$, let $\mathsf{P}_{r}:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ be the persistence module with $\mathsf{P}_{r}(s)=\mathbb{k}$ if $r\leq s$ and $\mathsf{P}_{r}(s)=0$ if $r\nleq s$, with all structure morphisms $\mathbb{k}\xrightarrow{\;\;\;}\mathbb{k}$ being the identity. Equivalently, one can define $\mathsf{P}_{r}$ to be $H_{0}(\{x\},\emptyset,\partial,f)$, with $f(x)=r$.
Notation 2.2.
If $I$ is a finite set and $f:I\xrightarrow{\;\;\;}\mathscr{P}$ is any function, we can consider the direct sum $M=\bigoplus_{i\in I}\mathsf{P}_{f(i)}:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$. When we need to work with elements of such a direct sum, we distinguish summands by writing $M=\bigoplus_{i\in I}\left(\mathsf{P}_{f(i)}\cdot\{i\}\right)$, so that $\left(\mathsf{P}_{f(i)}\cdot\{i\}\right)(r)=0$ if $r\ngeq f(i)$ and $\left(\mathsf{P}_{f(i)}\cdot\{i\}\right)(r)$ is equal to the free vector space generated by $\{i\}$, for $r\geq f(i)$.
Definition 2.3.
A persistence module $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ is projective of finite rank if there exists a function $\beta^{M}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$ of finite support such that $M\cong\bigoplus_{r\in\mathscr{P}^{n}}\mathsf{P}_{r}^{\beta^{M}(r)}$.
Note that, drawing inspiration from commutative algebra, projective persistence modules are sometimes also called free. The following result justifies the term projective used in Definition 2.3; see, e.g., [36, Section 3.1] for the usual notion of projective module.
Lemma 2.4.
Let $g:M\xrightarrow{\;\;\;}N$ be a surjection between $\mathscr{P}$-persistence modules, and let $h:P\xrightarrow{\;\;\;}N$ with $P$ projective of finite rank. There exists a morphism $h^{\prime}:P\xrightarrow{\;\;\;}M$ such that $g\circ h^{\prime}=h$.
Resolutions, presentations, and Betti tables
The next result is standard; see, e.g., [27, Proposition 6.24].
Lemma 2.5.
If $M$ is projective of finite rank, then there exists exactly one function (necessarily of finite support) $\beta^{M}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$ such that $M\cong\bigoplus_{r\in\mathscr{P}}\mathsf{P}_{r}^{\beta^{M}(r)}$.∎
Definition 2.6.
The Betti table of a persistence module $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ that is projective of finite rank is the function $\beta^{M}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$ of Lemma 2.5.
The following notation is sometimes convenient.
Notation 2.7.
If $r\in\mathscr{P}$, we let $\delta_{r}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$ be the function defined by $\delta_{r}(r)=1$ and $\delta_{r}(s)=0$ if $s\neq r$. In particular, $\delta_{r}=\beta^{\mathsf{P}_{r}}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$.
Definition 2.8.
Let $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ be a $\mathscr{P}$-persistence module. A finite projective cover of $M$ is a surjective morphism $P\xrightarrow{\;\;\;}M$ where $P$ is projective of finite rank, and $\sum_{r\in\mathscr{P}}\beta^{P}(r)\in\mathbb{N}$ is minimal.
Definition 2.9.
Let $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ be a $\mathscr{P}$-persistence module, and let $k\in\mathbb{N}$. A finite projective $k$-resolution (resp. minimal finite projective $k$-resolution) of $M$, denoted $C_{\bullet}\xrightarrow{\;\;\;}M$, is a sequence of morphisms $C_{k}\xrightarrow{\partial_{k}}C_{k-1}\xrightarrow{\partial_{k-1}}\cdots\xrightarrow{\partial_{2}}C_{1}\xrightarrow{\partial_{1}}C_{0}\xrightarrow{\partial_{0}}M$ satisfying
  • $C_{0}\xrightarrow{\;\;\;}M$ is surjective (resp. a projective cover);
  • $\partial_{i}\circ\partial_{i+1}=0$, for every $10\leq i\leq k-$ 1 (so that $\partial_{i+1}$ factors through $\operatorname{ker}\partial_{i}$);
  • $C_{i+1}\xrightarrow{\partial_{i+1}}\operatorname{ker}(\partial_{i})$ is surjective (resp. a projective cover), for every $10\leq i\leq k-$ 1;
  • $C_{i}$ is projective of finite rank, for every $0\leq i\leq k$.
In particular, a minimal finite projective 00-resolution is simply a projective cover.
Notation 2.10.
Since we only consider finite resolutions, we omit the word “finite” and simply say projective $k$-resolution. A (minimal) projective presentation is a (minimal) projective $1$-resolution.
Definition 2.11.
Let $k\in\mathbb{N}$. A persistence module $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ is finitely $k$-resolvable if it admits a finite projective $k$-resolution.
Definition 2.12.
Let $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ be finitely $k$-resolvable and let $0\leq i\leq k$. The $i$th Betti table of $M$ is the function $\beta^{M}_{i}:\mathscr{P}\xrightarrow{\;\;\;}\mathbb{N}$ (necessarily of finite support) defined as $\beta^{M}_{i}\coloneqq\beta^{C_{i}}$, where $C_{\bullet}\xrightarrow{\;\;\;}M$ is a minimal $k$-resolution of $M$.
Notation 2.13.
When convenient, we write $\beta_{i}(M)$ instead of $\beta_{i}^{M}$ for the Betti tables of a persistence module $M$.
The Betti tables of $M$ are also sometimes called the (multigraded) Betti numbers of $M$. The Betti tables of $M$, as defined in Definition 2.12, are independent of the choice of minimal presentation or resolution, thanks to the following standard result.
Lemma 2.14.
Let $k\in\mathbb{N}$. If $M:\mathscr{P}\xrightarrow{\;\;\;}\mathbf{Vec}$ is finitely $k$-resolvable, then it admits a minimal projective $k$-resolution $C_{\bullet}\xrightarrow{\;\;\;}M$. Moreover, any other projective $k$-resolution $C^{\prime}_{\bullet}\xrightarrow{\;\;\;}M$ has the property that $\beta^{C_{i}}\leq\beta^{C^{\prime}_{i}}$ for all $0\leq i\leq k$.
Example 2.15.
Consider the graph $G=(V,E,\partial)$ with $V=\{x,y\}$, $E=\{a,b\}$, and $\partial(a)=\partial(b)=(x,y)$. Consider the filtration $f:G\xrightarrow{\;\;\;}\mathbb{R}^{2}$ with $f(x)=f(y)=(0,0)$, $f(a)=(0,1)$, and $f(b)=(1,0)$. Then, a minimal resolution of $H_{0}(G,f)$ is given by
$0\xrightarrow{\;\;\;}\mathsf{P}_{(1,1)}\cdot\{\alpha\}\xrightarrow{\partial_{2}}\mathsf{P}_{(1,0)}\cdot\{a\}\oplus\mathsf{P}_{(0,1)}\cdot\{b\}\xrightarrow{\partial_{1}}\mathsf{P}_{(0,0)}\cdot\{x\}\oplus\mathsf{P}_{(0,0)}\cdot\{y\}\xrightarrow{\partial_{0}}H_{0}(G,f),$
where $\partial_{0}(\{x\})=[x]$, $\partial_{0}(\{y\})=[y]$, $\partial_{1}(\{a\})=\partial_{1}(\{b\})=\{y\}-\{x\}$, and $\partial_{2}(\{\alpha\})=\{a\}-\{b\}$. In particular, $\beta_{0}(H_{0}(G,f))=\delta_{(0,0)}+\delta_{(0,0)}$, $\beta_{1}(H_{0}(G,f))=\delta_{(1,0)}+\delta_{(0,1)}$, and $\beta_{2}(H_{0}(G,f))=\delta_{(1,1)}$. This can be easily checked by hand, but it also follows from Theorems 3.8 and 3.10, since $(G,f)$ is a minimal filtered graph, in the sense of Definition 3.2, which we give in the next section.
3 Theory
We start with a simple, standard result which says that a projective presentation of 00-dimensional homology is given by using the grades of vertices as generators, and the grades of edges as relations. Proofs can be found in Section A.2.
Lemma 3.1.
Let $\mathscr{P}$ be a poset, let $(G,f)$ be a $\mathscr{P}$-filtered graph, and consider the following morphism of projective modules
$\bigoplus_{e\in E}\mathsf{P}_{f(e)}\cdot\{e\}\;\;\;$ $\xrightarrow{\;\;\;\partial_{1}^{(G,f)}\;\;\;}\;\;\;\bigoplus_{v\in V}\mathsf{P}_{f(v)}\cdot\{v\}$
$\{e\}\;\;\;$$\xmapsto{\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;}\;\;\;\{e_{1}\}-\{e_{0}\}$
Then $H_{0}(G,f)\cong\operatorname{coker}\left(\partial_{1}^{(G,f)}\right)$ and $H_{1}(G,f)\cong\operatorname{ker}\left(\partial_{1}^{(G,f)}\right)$. In particular, $H_{0}(G,f)$ is a finitely presentable persistence module.
The point of minimal filtered graphs, which we now introduce, is that they make the presentation in Lemma 3.1 minimal (see Theorem 3.8).
Definition 3.2.
Let $(V,E,\partial,f)$ be a filtered graph and let $e\in E$.
  • The edge $e$ is collapsible if $e_{0}\neq e_{1}$ and $f(e)=f(e_{i})$ for some $i\in\{0,1\}$.
  • The edge $e$ is deletable if $[e_{0}]=[e_{1}]\in\pi_{0}(V,E\setminus\{e\},\partial,f)(f(e))$.
A filtered graph $(G,f)$ is
  • vertex-minimal if it does not contain any collapsible edges.
  • edge-minimal if it does not contain any deletable edges.
  • minimal if it is vertex-minimal and edge-minimal.
Example 3.3.
The graph of Example 2.15 is minimal, as neither of the two edges is collapsible or deletable. On the other hand, the graph of Fig. 1 is not minimal since $h_{1}$, $h_{2}$, $d_{1}$, $d_{2}$, and $d_{3}$ are collapsible, and $e_{3}$ is deletable.
As their name suggests, collapsible and deletable edges can be collapsed or deleted:
Construction 3.4.
Let $(G,f)=(V,E,\partial,f)$ be a filtered graph.
  • If $e\in E$ is collapsible, let $i\in\{0,1\}$ be such that $f(e)=f(e_{i})$. Define the simple collapse $G\downarrow e\coloneqq(V\setminus\{e_{i}\},E\setminus\{e\},\partial^{\prime})$, where $\partial^{\prime}=(\varphi\times\varphi)\circ\partial$ and $\varphi:V\xrightarrow{\;\;\;}V$ is given by $\varphi(v)=v$ if $v\neq e_{i}$ and $\varphi(v)=e_{1-i}$ if $v=e_{i}$.
  • If $e\in E$ is any edge, define the simple deletion $G\setminus e\coloneqq(V,E\setminus\{e\},\partial)$.
We now study the effect of collapsing collapsible edges, and of deleting deletable edges. We start with a useful observation, whose proof is straightforward.
Lemma 3.5.
Let $(G,f)$ be a filtered graph and let $e$ and $d$ be two different edges of $G$.
  1. 1.
    Assume that $e$ is collapsible. If $d$ is not collapsible (resp. deletable) in $(G,f)$, then $d$ is not collapsible (resp. deletable) in $(G\downarrow e,f)$.
  2. 2.
    If $d$ is not collapsible (resp. deletable) in $(G,f)$, then $d$ is not collapsible (resp. deletable) in $(G\setminus e,f)$.∎
Lemma 3.6.
If $(G,f)=(V,E,\partial,f)$ is a filtered graph and $e\in E$ is collapsible, then $H_{i}(G,f)\cong H_{i}(G\downarrow e,f)$ for $i\in\{0,1\}$.
Lemma 3.7.
If $(G,f)=(V,E,\partial,f)$ is a filtered graph and $e\in E$ is deletable, then $H_{0}(G,f)\cong H_{0}(G\setminus e,f)$ and $H_{1}(G,f)\cong H_{1}(G\setminus e,f)\oplus\mathsf{P}_{f(e)}$.
Next is a construction of a minimal presentation of $H_{0}$ of a minimal filtered graph.
Theorem 3.8.
Let $\mathscr{P}$ be a poset, let $(G,f)$ is a $\mathscr{P}$-filtered graph, and let $\partial_{1}^{(G,f)}:C_{1}\xrightarrow{\;\;\;}C_{0}$ be the morphism of Lemma 3.1.
  1. 1.
    If $(G,f)$ is vertex-minimal, then the map $C_{0}\xrightarrow{\;\;\;}\operatorname{coker}(\partial_{1}^{(G,f)})$ is a projective cover. In this case, $\beta_{0}\big{(}H_{0}(G,f)\big{)}=\sum_{v\in V}\delta_{f(v)}$.
  2. 2.
    If $(G,f)$ is a minimal filtered graph, then $C_{1}\xrightarrow{\partial_{1}^{(G,f)}}C_{0}\xrightarrow{\;\;\;}\operatorname{coker}(\partial_{1}^{(G,f)})$ is a minimal presentation. In this case, we also have $\beta_{1}\big{(}H_{0}(G,f)\big{)}=\sum_{e\in E}\delta_{f(e)}$.
We now focus on the case of $\mathbb{R}^{2}$-filtered graphs. If $(G,f)$ is an $\mathbb{R}^{2}$-filtered graph, let $f_{\mathbf{x}},f_{\mathbf{y}}:G\xrightarrow{\;\;\;}\mathbb{R}$ denote the first and second coordinates of $f$, respectively.
Definition 3.9.
Let $(V,E,\partial,f)$ be a minimal $\mathbb{R}^{2}$-filtered graph, and let $\prec$ be a total order on the set of edges $E$ that refines the lexicographic order induced by $f$ (i.e., $e\prec e^{\prime}$ implies that $f_{\mathbf{x}}(e)<f_{\mathbf{x}}(e^{\prime})$ or $f_{\mathbf{x}}(e)=f_{\mathbf{x}}(e^{\prime})$ and $f_{\mathbf{y}}(e)\leq f_{\mathbf{y}}(e^{\prime})$). An edge $d\in E$ is cycle-creating with respect to $\prec$ if $[d_{0}]=[d_{1}]\in\pi_{0}(V,\{e\in E:e\prec d\},\partial)$.
Thus, $d\in E$ is cycle-creating if there exists a list of edges $e^{1},\dots,e^{k}\in E$ with $e^{i}\prec d$ for $1\leq i\leq k$, and a list of signs $s^{1},\dots,s^{k}\in\{+,-\}$, such that $s^{1}e^{1},\dots,s^{k}e^{k}$ is a directed path from $d_{0}$ to $d_{1}$. Such a path $w=(s^{\bullet},e^{\bullet})$ is called a witness for $d$, and we denote $f(w)\coloneqq\left(\,f_{\mathbf{x}}(d)\,,\,\max_{1\leq i\leq k}f_{\mathbf{y}}(e^{i})\,\right)$. A minimal witness $w$ of a cycle-creating edge $d$ is one for which $f_{\mathbf{y}}(w)\in\mathbb{R}$ is as small as possible.
Since $\prec$ refines the lexicographic order, we have $f_{\mathbf{x}}(e^{i})\leq f_{\mathbf{x}}(d)$ for all $i$; thus, if $\max_{1\leq i\leq k}f_{\mathbf{y}}(e^{i})\leq f_{\mathbf{y}}(d)$, we would have that $[d_{0}]=[d_{1}]\in\pi_{0}(V,E\setminus\{d\},\partial,f)(f(d))$, which contradicts the fact that the filtered graph is minimal. This means that if $w$ is a witness for $d$, then $\max_{1\leq i\leq k}f_{\mathbf{y}}(e^{i})>f_{\mathbf{y}}(d)$, and thus $f(d)<f(w)$.
Theorem 3.10.
Let $(G,f)=(V,E,\partial,f)$ be a minimal $\mathbb{R}^{2}$-filtered graph, and $\prec$ a total order on $E$ refining the lexicographic order. For each $d\in E$ cycle-creating, let $w_{d}=(s_{d}^{\bullet},e_{d}^{\bullet})$ be a minimal witness wrt to $\prec$. The kernel of the morphism $\partial_{1}^{(G,f)}$ of Lemma 3.1 is given by:
$\bigoplus_{\begin{subarray}{c}d\in E\\ \text{cycle-creating}\end{subarray}}\mathsf{P}_{f(w_{d})}\cdot\{d\}\;\;\;$ $\xrightarrow{\;\;\kappa^{(G,f)}\;\;}\;\;\;\bigoplus_{e\in E}\mathsf{P}_{f(e)}\cdot\{e\}$
$\{d\}\;\;\;$$\xmapsto{\;\;\;\;\;\;\;\;\;\;\;\;\;\;}\;\;\;\{d\}-\big{(}s_{d}^{1}\{e_{d}^{1}\}+\cdots+s_{d}^{k}\{e_{d}^{k}\}\big{)}.$
It follows that
  • $\beta_{0}(H_{0}(G,f))=\sum_{v\in V}\delta_{f(v)}$;
  • $\beta_{1}(H_{0}(G,f))=\sum_{e\in E}\delta_{f(e)}$;
  • $\beta_{0}(H_{1}(G,f))=\beta_{2}(H_{0}(G,f))=\sum_{d\in E,\text{cycle-creating}}\delta_{f(w_{d})}$;
  • $\beta_{i}(H_{0}(G,f))=0$ for $i\geq 3$, and $\beta_{i}(H_{1}(G,f))=0$ for $i\geq 1$.
4 Algorithms
Outline
We first introduce some technical notions and overview the algorithms.
Two $\mathscr{P}$-filtered graph $(G,f)$ and $(G^{\prime},f^{\prime})$ are homology-equivalent (over $\mathbb{k}$) if $H_{0}(G,f;\mathbb{k})\cong H_{0}(G^{\prime},f^{\prime};\mathbb{k})$ and $H_{1}(G,f;\mathbb{k})\cong H_{1}(G^{\prime},f^{\prime};\mathbb{k})$.
Algorithm 2 reduces the input filtered graph to a vertex-minimal filtered graph by collapsing edges; it relies on Algorithm 1, which collapses local collapsible edges. A local collapsible edge of $(G,f)$ is an edge $e$ such that $e_{0}\neq e_{1}$ and $f(e)=f(e_{0})=f(e_{1})$. Algorithm 2 then identifies minimal vertices, that is, vertices with no adjacent collapsible edge decreasing the grade, and runs a depth-first search from each of these to build and collapse a tree of collapsible edges.
Algorithm 3 first calls Algorithm 2 to perform all collapses and then deletes deletable edges until there are no more deletable edges. It then builds the presentation of Lemma 3.1.
Algorithm 4 first calls Algorithm 2 to perform all collapses, and then goes through all edges, with respect to the lex order on their grades, and identifies deletable edges, non-deletable edges, and cycle-creating edges.
Dynamic dendrograms
Since this is used in Algorithm 4, we now introduce the dynamic dendrogram data structure, which, informally, represents a dendrogram where elements merge as time goes from $\infty$- ∞ to $\infty$, and is dynamic in that one is allowed to change the dendrogram by making two elements merge earlier.
Definition 4.1.
Let $(G,f)=(V,E,\partial,f)$ be a [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-filtered graph. A dynamic dendrogram $D$ for $(G,f)$ is a data structure supporting the following operations:
  • If $v,w\in V$, the operation $D.\mathsf{time\_of\_merge}(v,w)$ returns the smallest $r\in[-\infty,\infty)$ such that $[v]=[w]\in\pi_{0}(G,f)(r)$, or $\infty$ if $[v]\neq[w]\in\pi_{0}(G,f)(r)$ for all $r\in[-\infty,\infty)$.
  • If $v,w\in V$ and $t\in[-\infty,\infty)$, the operation $D.\mathsf{merge\_at\_time}(v,w,t)$ modifies the dendrogram so that it is a dynamic dendrogram for $(V,E\sqcup\{e\},\partial^{\prime},f^{\prime})$, with $f^{\prime}|_{E}=f$, $\partial^{\prime}|_{E}=\partial$, $f^{\prime}(e)=t$, and $\partial^{\prime}(e)=(v,w)$.
A dynamic dendrogram $D$ can easily and efficiently be implemented using a mergeable tree $T$ in the sense of [21]. These trees store heap-ordered forests, i.e., a collection of rooted labeled trees, where the labels decrease (or in our case increase) along paths to the root. The data structures support a wealth of operations for dynamic updates. We need only three of them: $\mathsf{insert}(v,t)$ inserts node $v$ with label $t$ into the forest; $\mathsf{nca}(v,w)$ finds the nearest common ancestor of two nodes $v$ and $w$; $\mathsf{merge}(v,w)$ merges the paths of $v$ and $w$ to their root(s) while preserving the heap order. We use these operations to implement dynamic dendrograms as follows:
  • To implement $D.\mathsf{time\_of\_merge}(v,w)$, return the label of $T.\mathsf{nca}(v,w)$.
  • To implement $D.\mathsf{merge\_at\_time}(v,w,t)$, let $h$ be a new vertex not already in the mergeable tree $T$, do $T.\mathsf{insert}(h,t)$, then $T.\mathsf{merge}(v,h)$, and $T.\mathsf{merge}(w,h)$.
Presentation matrices
In the algorithms, to represent a Betti table $\beta_{i}(M)$ we use a list of elements of the indexing poset. If we have represented the 00th and $1$st Betti tables of $M$ with lists $\beta_{0}$ and $\beta_{1}$, we represent a minimal presentation for $M$ with a sparse matrix using coordinate format, that is, with a list of triples $(i,j,v)$ representing the fact that $v$ is the entry at row $i$ and column $j$, and where $i$ represents the $i$th element of $\beta_{0}$ and $j$ represents the $j$th element of $\beta_{1}$.
Input: Filtered graph $(G,f)=(V,E,\partial,f)$
Output: Filtered graph $(G^{\prime},f)$ homology-equivalent to $(G,f)$ and without local collapsible edges
1 Initialize dictionary $\varphi$ with identity map, $\varphi[v]=v$ for $v\in V$
2 Initialize empty set ${\sf visited}$
3 Initialize $E^{\prime}\leftarrow E$
4 for $v\in V$ do $\triangleright$ Run depth-first search from $v$ on local edges
5       Initialize ${\sf stack}$ with $(v,\emptyset,v)$
6       while ${\sf stack}$ is not empty do
7             $(v,e,u)\leftarrow{\sf stack}.{\bf pop}()$
8             if $u\notin{\sf visited}$ then
9                   add $u$ to ${\sf visited}$
10                   if $e\neq\emptyset$ then
11                         $\varphi[u]\leftarrow v$
12                         remove $e$ from $E^{\prime}$
13
14                  for $e\in E$ with $\{e_{0},e_{1}\}=\{u,x\}$ and $f(e)=f(x)=f(u)$ do $\triangleright$ local collapsible
15                         push $(v,e,x)$ onto ${\sf stack}$
16
[…]
return $(G^{\prime},f)=\left(V^{\prime},E^{\prime},(\varphi\times\varphi)\circ\partial,f\right)$
Algorithm 1 Collapse local collapsible edges
Input: Filtered graph $(G,f)=(V,E,\partial,f)$
Output: Vertex-minimal filtered graph $(G^{\prime},f)$ homology-equivalent to $(G,f)$, and $\beta_{0}(H_{0}(G,f))$
[…]
2 Initialize dictionary $\varphi$ with identity map, $\varphi[v]=v$ for $v\in V_{i}$
3 Initialize empty set ${\sf visited}$
4 Initialize $E^{\prime}\leftarrow E_{i}$
5 for $v$ in $V_{i}$ do $\triangleright$ Run depth-first search from minimal vertices
6       if $\exists$ edge $e\in E_{i}$ with $\{e_{0},e_{1}\}=\{u,v\}$ and $f(e)=f(v)>f(u)$ then $\triangleright$ $v$ is not minimal
7             continue
8
9      Initialize ${\sf stack}$ with $(v,\emptyset,v)$
10       while ${\sf stack}$ is not empty do
11             $v,e,u\leftarrow{\sf stack}.{\bf pop}()$
12             if $u\notin{\sf visited}$ then
13                   add $u$ to ${\sf visited}$
14                   if $e\neq\emptyset$ then
15                         $\varphi[u]\leftarrow v$
16                         remove $e$ from $E^{\prime}$
17
18                  for every edge $e\in E_{i}$ with $\{e_{0},e_{1}\}=\{u,x\}$ with $f(e)=f(x)>f(u)$ do
19                         push $(v,e,x)$ onto ${\sf stack}$
20
[…]
22 return $(G^{\prime},f)=\left(V^{\prime},E^{\prime},(\varphi\times\varphi)\circ\partial_{i},f\right)$ and $\beta_{0}=[f(v):v\in V^{\prime}]$
Algorithm 2 Collapse to vertex-minimal filtered graph
Input: Filtered graph $(G,f)=(V,E,\partial,f)$
Output: Minimal presentation of $H_{0}(G,f;\mathbb{k})$
1 Initialize $\beta_{0},\beta_{1}$ with empty lists
2 Initialize empty sparse matrix $\mathcal{M}$ and empty dictionary $\mathsf{row\_idx}$
[…]
4 for $e\in E^{\prime}$  do $\triangleright$ Check each edge, and delete it if it is deletable
5       Define set of vertices $V_{e}\leftarrow\{v\in V^{\prime}:f(v)\leq f(e)\}$
6       Define set of edges $E_{e}\leftarrow\{d\in E^{\prime}:f(d)\leq f(e)\}\setminus e$
7       Run breadth-first search on $(V_{e},E_{e},\partial^{\prime})$ starting from $e_{0}$
8       if $e_{1}$ is reachable from $e_{0}$ then $\triangleright$ $[e_{0}]=[e_{1}]\in\pi_{0}(V_{e},E_{e},\partial^{\prime})$ so $e$ is deletable
9             $E^{\prime}\leftarrow E^{\prime}\setminus\{e\}$
10
11for $v\in V^{\prime}$ do
12       $\beta_{0}.\mathsf{append}(f(v))$
13       $\mathsf{row\_idx}[v]\leftarrow|\beta_{0}|$
14
15for $e\in E^{\prime}$ do $\triangleright$ The morphism $\partial_{1}^{(V^{\prime},E^{\prime},\partial^{\prime})}$ of Lemma 3.1
16       $\beta_{1}.\mathsf{append}(f(e))$
17       .$\mathcal{M}.\mathsf{append}(\,\mathsf{row\_idx}[e_{0}]\,,\,|\beta_{1}|\,,\,-1\,)$
[…]
19
return $\beta_{0},\beta_{1},\mathcal{M}$
Algorithm 3 Minimal presentation of $\mathscr{P}$-filtered graph
Input: Filtered graph $(G,f)=(V,E,\partial,f)$
Output: Betti tables $\beta_{0},\beta_{1},\beta_{2}$ and minimal presentation of $H_{0}(G,f;\mathbb{k})$, and $\beta_{0}^{1}\coloneqq\beta_{0}(H_{1}(G,f;\mathbb{k}))$
1 Initialize $\beta_{0},\beta_{1},\beta_{2},\beta^{1}_{0}$ with empty lists
2 Initialize empty sparse matrix $\mathcal{M}$ and empty dictionary $\mathsf{row\_idx}$
[…]
4 Let $D$ be a dynamic dendrogram on $(V^{\prime},\emptyset,\partial,g)$, with $g(v)=-\infty$ for all $v\in V^{\prime}$
5 for $(x,y)\in\mathbb{R}^{2}$ s.t. $f^{-1}(x,y)\neq\emptyset$ in lex order do $\triangleright$ Visit grades lexicographically
6       for $v\in V^{\prime}\text{ with }f(v)=(x,y)$  do $\triangleright$ All vertices belong to the projective cover
7             $\beta_{0}.\mathsf{append}(\,(x,y)\,)$
8             $\mathsf{row\_idx}[v]\leftarrow|\beta_{0}|$
9
10      for $e\in E^{\prime}\text{ with }f(e)=(x,y)$  do
11             $s\leftarrow D.\mathsf{time\_of\_merge}(e_{0},e_{1})$
12             if $s\leq y$  then $\triangleright$ The edge is deletable, so it only affects $H_{1}$
13                   $\beta^{1}_{0}.\mathsf{append}(\,(x,y)\,)$
14
15            else  $\triangleright$ Edge is not deletable, so belongs to relations in resolution
16                   $D.\mathsf{merge\_at\_time}(e_{0},e_{1},y)$
17                   $\beta_{1}.\mathsf{append}(\,(x,y)\,)$
18                   .$\mathcal{M}.\mathsf{append}(\,\mathsf{row\_idx}[e_{0}]\,,\,|\beta_{1}|\,,\,-1\,)$
[…]
20                   if $s<\infty$ then $\triangleright$ The edge is cycle-creating
21                         $\beta_{2}.\mathsf{append}(\,(x,s)\,)$
22                         $\beta^{1}_{0}.\mathsf{append}(\,(x,s)\,)$
23
return $\beta_{0},\beta_{1},\beta_{2},\beta^{1}_{0},\mathcal{M}$
Algorithm 4 Betti tables and minimal presentation of $\mathbb{R}^{2}$-filtered graph
Complexity and correctness
We conclude the paper by using the theoretical results of Section 3 to prove the main results in the introduction. Proofs for the results in this section are in Section A.3.
We start by with a convenient lemma, which gives conditions under which one can collapse an entire subgraph and produce a homology-equivalent filtered graph. The following definition describes the type of subgraph that can be collapsed.
Definition 4.2.
Let $(G,f)=(V,E,\partial,f)$ be a filtered graph. A subset $E^{\prime}\subseteq E$ is a monotonic forest of $(G,f)$ if:
  1. 1.
    The subgraph of $G$ spanned by the edges in $E^{\prime}$ is a forest.
  2. 2.
    For every vertex $v$ in the forest, there exists at most one edge $e\in E^{\prime}$ such that $\{e_{0},e_{1}\}=\{v,w\}$ and $f(w)<f(v)$.
Lemma 4.3.
Let $(G,f)=(V,E,\partial,f)$ be a filtered graph, and let $E^{\prime}\subseteq E$ be a set of collapsible edges, which forms a monotonic forest. If $e\in E^{\prime}$, then all the edges in $E^{\prime}\setminus\{e\}$ are collapsible in $(G\downarrow e,f)$. In particular, the whole forest $E^{\prime}$ can be collapsed (in any order) to obtain a filtered graph that is homology-equivalent to $(G,f)$.
Lemma 4.4.
Let $(G,f)$ be a $\mathscr{P}$-filtered graph. Algorithm 1 outputs a filtered graph homology-equivalent to $(G,f)$ and without local collapsible edges in time $O(|G|)$.
Proposition 4.5.
Let $(G,f)$ be a $\mathscr{P}$-filtered graph. Algorithm 2 outputs a vertex-minimal filtered graph homology-equivalent to $(G,f)$ and $\beta_{0}(H_{0}(G,f))$ in $O(|G|)$ time.
Proposition 4.6.
Let $(G,f)$ be a finite $\mathscr{P}$-filtered graph. Algorithm 3 outputs a minimal presentation of $H_{0}(G,f)$ in $O(\,|G|^{2}\,)$ time.
With these results, we prove Theorem A and Theorem B in Section A.4.
Appendix A Proofs
A.1 Proofs from Section 2
Proof A.1 (Proof of Lemma 2.4).
Let $P=\bigoplus_{i\in I}\mathsf{P}_{f(i)}\cdot\{i\}$. To define a morphism $P\xrightarrow{\;\;\;}M$ it is sufficient to say, for each $j\in J$, to which element of $M(f(j))$ the element $\{j\}\in P(f(j))$ gets mapped to. Since $g$ is a surjection, let us choose $m_{j}\in M(f(j))$ in the preimage of $h_{f(j)}(\{j\})\in N(f(j))$, for each $j\in J$. Let $h^{\prime}:P\xrightarrow{\;\;\;}M$ be defined by mapping $\{j\}\in P(f(j))$ to $m_{j}\in M(f(j))$. It is clear that this morphism satisfies $g\circ h^{\prime}=h$.
Lemma A.2.
Let $M$ be a $\mathscr{P}$-persistence module and let $g:Q=\bigoplus_{i\in I}\mathsf{P}_{f(i)}\cdot\{i\}\xrightarrow{\;\;\;}M$ be a finite projective cover. For every $j\in I$, we have that $g_{f(j)}(\{j\})\neq 0\in M(f(j))$, and $\{j\}\in Q(f(j))$ is the only preimage of $g_{f(j)}(\{j\})$.
Proof A.3.
If $g_{f(j)}(\{j\})=0\in M(f(j))$, then the restriction of $g$ given by $\bigoplus_{i\in I\setminus\{j\}}\mathsf{P}_{f(i)}\cdot\{i\}\xrightarrow{\;\;\;}M$ would be surjective, and thus $\sum_{r\in\mathscr{P}}\beta^{Q}(r)$ would not be minimal, so $Q\xrightarrow{\;\;\;}M$ would not be a projective cover.
Similarly, if there exists another element of $Q(f(j))$ mapping to $g_{f(j)}(\{j\})$, then this element would be linearly independent of $\{j\}$, and again the restriction of $g$ given by $\bigoplus_{i\in I\setminus\{j\}}\mathsf{P}_{f(i)}\cdot\{i\}\xrightarrow{\;\;\;}M$ would be surjective, and $Q\xrightarrow{\;\;\;}M$ would not be a projective cover.
Proof A.4 (Proof of Lemma 2.14).
We start by proving that any finitely 00-resolvable $M$ admits a projective cover. By definition, there exists a surjection $P\xrightarrow{\;\;\;}M$ from a projective of finite rank $P$. If $\sum_{r\in\mathscr{P}}\beta^{P}(r)\in\mathbb{N}$ is not minimal, we can remove at least one direct summand of $P$ to get $P^{\prime}\xrightarrow{\;\;\;}M$ still a surjection. Proceeding in this way, we necessarily end up with a projective cover of $M$.
Now, a straightforward inductive argument shows that any $k$-resolvable module admits a minimal projective $k$-resolution.
We now prove that, if $g:P\xrightarrow{\;\;\;}N$ is a projective cover and $h:Q\xrightarrow{\;\;\;}N$ is a surjection from a finitely presentable projective, then $\beta^{P}\leq\beta^{Q}$. The inequality in the statement of the result involving Betti tables follows from this.
By Lemma 2.4, there exist morphisms $g^{\prime}:P\xrightarrow{\;\;\;}Q$ and $h^{\prime}:Q\xrightarrow{\;\;\;}P$ such that $h\circ g^{\prime}=g$ and $g\circ h^{\prime}=h$. We claim that $h^{\prime}\circ g^{\prime}:P\xrightarrow{\;\;\;}P$ is the identity; this is sufficient since then $Q\cong P\oplus P^{\prime}$ for some finitely presentable projective $Q$, and thus $\beta^{P}\leq\beta^{P}+\beta^{P^{\prime}}=\beta^{Q}$.
Let $P=\bigoplus_{i\in I}\mathsf{P}_{f(i)}\cdot\{i\}$. Let $j\in J$. By Lemma A.2, we have that $(h^{\prime}\circ g^{\prime})_{f(j)}(\{j\})=\{j\}\in P(f(j))$, since $g=g\circ(h^{\prime}\circ g^{\prime})$ and $\{j\}$ is the unique element of $P(f(j))$ mapping to $g_{f(j)}(\{j\})\in M(f(j))$. Since the elements $\{j\}\in P(f(j))$ generate $P$, this implies that $h^{\prime}\circ g^{\prime}$ is the identity.
A.2 Proofs from Section 3
Proof A.5 (Proof of Lemma 3.1).
It is straightforward to see that the evaluation of the natural transformation $\partial_{1}^{(G,f)}$ at $r\in\mathscr{P}$ coincides with morphism Equation 1, where the graph is ($\left(\{v\in V:f(v)\leq r\},\{e\in E:f(e)\leq r\},\partial\right)$. Then $H_{0}(G,f)$ is finitely presentable by definition.
Proof A.6 (Proof of Lemma 3.6).
This follows at once from the homotopy-invariance property of homology; however, in order to make the exposition self-contained, let us give an algebraic proof, using the definition of homology for graphs given in Section 2.
Let $e\in E$ be collapsible, and assume that $f(e)=f(e_{0})$; the case $f(e)=f(e_{1})$ is analogous. Consider the following diagram
${\displaystyle\left(\bigoplus_{d\in E\setminus\{e\}}\mathsf{P}_{f(d)}\cdot\{d\}\right)\oplus\left(\mathsf{P}_{f(e)}\cdot\{e\}\right)}$${\displaystyle\left(\bigoplus_{v\in V\setminus\{e_{i}\}}\mathsf{P}_{f(v)}\cdot\{v\}\right)\oplus\left(\mathsf{P}_{f(e_{i})}\cdot\{e_{0}\}\right)}$${\displaystyle\bigoplus_{d\in E}\mathsf{P}_{f(d)}\cdot\{d\}}$${\displaystyle\bigoplus_{v\in V}\mathsf{P}_{f(v)}\cdot\{v\}\;.}$$\left[\begin{array}[]{c|c}\partial_{1}^{(G\downarrow e,f)}&0\\ \hline\cr 0&1\end{array}\right]$ $\partial_{1}^{(G,f)}$[$\left[\begin{array}[]{c}\{d\}\mapsto\begin{cases}\{d\}&\text{if $d_{1}\neq e_{0}$}\\ \{d\}+\{e\}&\text{else}\end{cases}\\ \hline\cr\{e\}\mapsto\{e\}\end{array}\right]^{T}$ [$\left[\begin{array}[]{c}\{v\}\mapsto\{v\}\\ \hline\cr\{e_{0}\}\mapsto\{e_{1}\}-\{e_{0}\}\end{array}\right]^{T}$
It is clear that the cokernel (resp. kernel) of the top horizontal morphism is isomorphic to the cokernel (resp. kernel) of $\partial_{1}^{(G\downarrow e,f)}$, which in turn is isomorphic to the 00th (resp. $1$st) persistent homology of $(G\downarrow e,f)$, by Lemma 3.1. So it is sufficient to prove that the diagram is commutative, and that the vertical morphisms are isomorphisms. The commutativity of the diagram is a straightforward check, using the definition of $G\downarrow e$ (3.4) in the case where $d\in E$ is such that $d_{1}=e_{0}$. The fact that the vertical morphisms are isomorphisms is proven by a simple Gaussian elimination using the matrix representation of the vertical morphisms.
Lemma A.7.
If $(G,f)=(V,E,\partial,f)$ is a filtered graph and $e\in E$ is deletable, then the morphism of persistent sets $\pi_{0}(G\setminus e,f)\xrightarrow{\;\;\;}\pi_{0}(G,f)$ induced by the inclusion $G\setminus e\xrightarrow{\;\;\;}G$ is an isomorphism.
Proof A.8.
The morphism $\pi_{0}(G\setminus e,f)\xrightarrow{\;\;\;}\pi_{0}(G,f)$ is clearly surjective, since $G\setminus e\xrightarrow{\;\;\;}G$ is surjective on vertices. Thus, it is sufficient to prove that, for all $r\in\mathscr{P}$ and all $x,y\in V$ with $f(x),f(y)\geq r$, we have that $[x]=[y]\in\pi_{0}(G,f)(r)$ implies $[x]=[y]\in\pi_{0}(G\setminus e,f)(r)$. So assume that $[x]=[y]\in\pi_{0}(G,f)(r)$, and let $e^{1},\dots,e^{k}$ be a path between $x$ and $y$ in $(G,f)_{r}$. If the path does not contain the edge $e$, then it is also a path between $x$ and $y$ in $(G\setminus e,f)_{r}$, and we are done. If the path does contain the edge $e$, then, in particular, $r\geq f(e)$. Since $e$ is deletable, there exists a path $d^{1},\dots,d^{\ell}$ between the vertices $e_{0}$ and $e_{1}$ in $(G\setminus e,f)_{f(r)}$. Now, take the path $e^{1},\dots,e^{k}$ and replace every occurrence of the edge $e$ with the path $d^{1},\dots,d^{\ell}$. This yields a path between $x$ and $y$ in $(G\setminus,f)_{r}$, so that $[x]=[y]\in\pi_{0}(G\setminus e,f)_{r}$, as required.
Proof A.9 (Proof of Lemma 3.7).
Lemma A.7, together with Lemma 2.1, implies that the morphism $H_{0}(G,f)\xrightarrow{\;\;\;}H_{0}(G\setminus e,f)$ induced by the inclusion of graphs is an isomorphism of persistence modules.
We now prove that $H_{1}(G,f)\cong H_{1}(G\setminus e,f)\oplus\mathsf{P}_{f(e)}$. Consider the following diagram:
[…]
The first and last modules in the second and third column are, respectively, a kernel and a cokernel, by Lemma 3.1, and the second and third row are clearly exact. Thus, the snake lemma [36, Corollary 6.12] implies the existence of the dashed morphism and the exactness of the sequence of modules which includes the dashed morphism. Moreover, the statement about $H_{0}$ proven in the first part of the proof implies the isomorphism in the bottom row. We thus have that $00\xrightarrow{\;\;\;}H_{1}(G\setminus e,f)\xrightarrow{\;\;\;}H_{1}(G,f)\xrightarrow{\;\;\;}\mathsf{P}_{f(e)}\xrightarrow{\;\;\;}$ 0 is a short exact sequence, which necessarily splits since $\mathsf{P}_{f(e)}$ is projective [36, Proposition 3.3]. The result follows.
Proof A.10 (Proof of Theorem 3.8).
The claim about the Betti tables follows directly from the claim that the presentation $\partial_{1}^{(G,f)}$ is minimal, so let us prove this claim.
We start by proving (1). Let $G=(V,E,\partial)$ be vertex-minimal. We first prove that $h:Q\coloneqq\bigoplus_{v\in V}\mathsf{P}_{f(v)}\cdot\{v\}\xrightarrow{\;\;\;}\operatorname{coker}{\partial_{1}^{(G,f)}}\cong H_{0}(G,f)\cong\langle\pi_{0}(G,f)(f(v))\rangle_{\mathbb{k}}$ is a projective cover, where in the last isomorphism we used Lemma 2.1. This morphism maps $\{v\}\in Q(f(v))$ to $[v]\in\langle\pi_{0}(G,f)(f(v))\rangle_{\mathbb{k}}$, for every $v\in V$. If the morphism were not a projective cover, then there would exist $v^{\prime}\in V$ such that the restriction $h^{\prime}:\bigoplus_{v\in V\setminus\{v^{\prime}\}}\mathsf{P}_{f(v)}\cdot\{v\}\xrightarrow{\;\;\;}\operatorname{coker}{\partial_{1}^{(G,f)}}\cong H_{0}(G,f)\cong\langle\pi_{0}(G,f)(f(v))\rangle_{\mathbb{k}}$ would still be surjective. This is not possible: $(G,f)$ is vertex-minimal, so $[v]\neq[v^{\prime}]\in\pi_{0}(G,f)(f(v^{\prime}))$ for every $v\in V\setminus\{v^{\prime}\}$, and thus $h^{\prime}$ would not be surjective, since $[v^{\prime}]\in\langle\pi_{0}(G,f)(f(v))\rangle_{\mathbb{k}}$ would not be in its image.
The claim about $\beta_{0}$ follows at once from the definition of Betti table (Definition 2.12).
We now prove (2). Let $G=(V,E,\partial)$ be a minimal filtered graph. We prove that $g:\bigoplus_{e\in E}\mathsf{P}_{f(e)}\cdot\{e\}\xrightarrow{\;\;\;}\operatorname{ker}(h)$ is a projective cover. If it were not, then there would exist $e^{\prime}\in E$ such that the restriction $g^{\prime}:P\coloneqq\bigoplus_{e\in E\setminus\{e^{\prime}\}}\mathsf{P}_{f(e)}\cdot\{e\}\xrightarrow{\;\;\;}\operatorname{ker}(h)$ of $g$ would still be surjective. Let $G^{\prime}=(V,E\setminus\{e\},\partial)$. In particular, the last two vertical morphisms in the following commutative diagram would be isomorphisms
[…]
This is not possible, since the last vertical morphism maps $[e^{\prime}_{0}],[e^{\prime}_{1}]\in\langle\pi_{0}(G^{\prime},f)(f(e^{\prime}))\rangle_{\mathbb{k}}$ to $[e^{\prime}_{0}],[e^{\prime}_{1}]\in\langle\pi_{0}(G,f)(f(e^{\prime}))\rangle_{\mathbb{k}}$, and, $[e^{\prime}_{0}]=[e^{\prime}_{1}]\in\langle\pi_{0}(G,f)(f(e^{\prime}))\rangle_{\mathbb{k}}$, since $e^{\prime}$ is an edge of $G$, while $[e^{\prime}_{0}]\neq[e^{\prime}_{1}]\in\langle\pi_{0}(G^{\prime},f)(f(e^{\prime}))\rangle_{\mathbb{k}}$, by the fact that $(G,f)$ is edge-minimal.
As for the case of (1), the claim about $\beta_{1}$ follows from the definition of Betti table.
Proof A.11 (Proof of Theorem 3.10).
The claim about the Betti tables follows at once from the first statement of the theorem together with Theorem 3.8. Note in particular that we need not prove that $\kappa^{(G,f)}$ and $\partial_{1}^{(G,f)}$ form a minimal resolution, since we know that $\partial_{1}^{(G,f)}$ is a minimal presentation. So let us prove the claim about the kernel.
We proceed by induction on the number of edges in the graph. Let $h\in E$ be the maximum of $E$ with respect to the total order $\prec$. The filtered graph $(G^{\prime},f)\coloneqq(G\setminus h,f)$ is minimal, by Lemma 3.5(2), so the result holds for this graph.
We consider the following diagram, obtained by adding the edge $h$ to $(G^{\prime},f)$:
00{0}${\displaystyle\bigoplus_{\text{\emph{cc}}\,d\in E\setminus\{h\}}\mathsf{P}_{f(w_{d})}\cdot\{d\}}$${\displaystyle\bigoplus_{e\in E\setminus\{h\}}\mathsf{P}_{f(e)}\cdot\{e\}}$${\displaystyle\bigoplus_{v\in V}\mathsf{P}_{f(v)}\cdot\{v\}}$${H_{0}(G^{\prime},f)}$00{0}00{0}${\displaystyle\bigoplus_{\text{\emph{cc}}\,d\in E}\mathsf{P}_{f(w_{d})}\cdot\{d\}}$${\displaystyle\bigoplus_{e\in E}\mathsf{P}_{f(e)}\cdot\{e\}}$${\displaystyle\bigoplus_{v\in V}\mathsf{P}_{f(v)}\cdot\{v\}}$${H_{0}(G,f)}$${0\,,}$$\kappa^{(G^{\prime},f)}$$\kappa^{(G,f)}$$\partial_{1}^{(G^{\prime},f)}$$\partial_{1}^{(G,f)}$
where cc stands for “cycle-creating”. Since $h$ is the maximum according to $\prec$, the cycle-creating edges of $(G^{\prime},f)$ with respect to $\prec$ are simply the cycle-creating edges of $(G,f)$, possibly minus $h$ in the case where $h$ is cycle-creating.
The top row of the diagram is exact by inductive hypothesis, and we need to prove that the bottom row is exact as well.
In order to prove this, we evaluate the persistence modules and morphisms of the diagram at each $(x,y)\in\mathbb{R}^{2}$, and show that the bottom row is exact. We consider three cases:
  1. 1.
    $(x,y)\not\geq f(h)$;
  2. 2.
    $(x,y)\geq f(h)$, and either $h$ is not cycle-creating or $(x,y)\not\geq f(w_{h})$;
  3. 3.
    $(x,y)\geq f(w_{h})$, which is only considered if $h$ is cycle-creating.
Case (1) is immediate, since the top and bottom row are equal when evaluated at $(x,y)$, as $h$ is not an edge of $(G,f)_{(x,y)}$.
Case (2) is slightly more interesting. In this case, the difference between the two rows, when evaluated at $(x,y)$, is that the edge $h$ has appeared in the second row, and it is connecting the connected components of $h_{0}$ and $h_{1}$, which would otherwise be distinct. This is because either $h$ is not cycle-creating, or $(x,y)\not\geq f(w_{h})$, so that $y<f_{\mathbf{y}}(w_{h})$, which is the smallest grade at which $h_{0}$ and $h_{1}$ become connected in $\{(x^{\prime},y^{\prime})\in\mathbb{R}^{2}:(x^{\prime},y^{\prime})\geq f(h)\}$, by the fact that $w_{h}$ is a minimal witness. In going from the top row to the bottom row, both evaluated at $(x,y)$, the dimension of $H_{0}$ is dropping by one, and the dimension of the vector space generated by edges is going up by one. Since, in the bottom row, the generator $\{h\}$ is mapping to $\{h_{1}\}-\{h_{0}\}$, which is non-zero in $H_{0}(G^{\prime},f)(x,y)$, this case follows from the following basic claim from linear algebra, whose proof is straightforward.
[…]
Finally, we consider case (3). Thus, in going from the first to the second row, when evaluated at $(x,y)$, the dimension of the vector space generated by edges is going up by one, and the dimension of the space generated by vertices remains the same. In this case, in $H_{0}(G^{\prime},f)(x,y)$, the elements $[h_{0}]$ and $[h_{1}]$ are already equal, since there is a path between $h_{0}$ and $h_{1}$, namely, the witness $w_{h}$; thus, the dimension of $\operatorname{ker}\partial_{1}(x,y)$ is going up by one. In order to account for this, $\kappa^{(G,f)}$ is mapping $\{h\}$ to $\{h\}-(s^{1}_{h}\{e^{1}_{d}\}+\cdots+s^{k}_{h}\{e^{k}_{d}\})$, which is clearly non-zero, and in the kernel of $\partial_{1}(x,y)$. This concludes the proof.
A.3 Proofs from Section 4
Proof A.12 (Proof of Lemma 4.3).
Collapsing a single collapsible edge produces a homology-equivalent filtered graph, by Lemma 3.6. So we need to prove that collapsing an edge of the monotonic forest $E^{\prime}$ consisting of collapsible edges does not affect the collapsibility of other edges in $E^{\prime}$.
Let $e,d\in E^{\prime}$. We prove that $d$ is collapsible in $(G\downarrow e,f)$. Without loss of generality, we assume that $f(d)=f(d_{1})$. Without loss of generality, we also assume that $f(e)=f(e_{1})$ and that the simple collapse deletes $e_{1}$ and $e$, and makes every edge adjacent to $e_{1}$ now be adjacent to $e_{0}$.
Let $G^{\prime}=(V^{\prime},E\setminus\{e\},\partial^{\prime})=G\downarrow e$. For notational convenience, let $(d^{\prime}_{0},d^{\prime}_{1})=\partial^{\prime}(d)$. We prove that $d^{\prime}_{0}\neq d^{\prime}_{1}$ and $f(d^{\prime}_{0})=f(d)$ or $f(d^{\prime}_{1})=f(d)$, which implies that $d$ is collapsible in $(G\downarrow d,f)$, by definition. The fact that $d^{\prime}_{0}\neq d^{\prime}_{1}$ in $G\downarrow e$ is clear, since $d_{0}\neq d_{1}$, so $d^{\prime}_{0}=d^{\prime}_{1}$ would mean that $\{e_{0},e_{1}\}=\{d_{0},d_{1}\}$, which is not possible because $E^{\prime}$ is a forest and thus contains no loops.
We prove $f(d^{\prime}_{0})=f(d)$ or $f(d^{\prime}_{1})=f(d)$ by considering two cases. The first case is when we have $d_{1}\neq e_{1}$. In this case $d^{\prime}_{1}=d_{1}$, so $f(d^{\prime}_{1})=f(d_{1})=f(d)$. The second case is when $d_{1}=e_{1}$. In this case, we cannot have $f(d_{0})<f(d)$, as this would violate condition (2) of Definition 4.2 for the vertex $v=e_{1}$, so we must have $f(d_{0})=f(d)$. But since $d_{0}\neq d_{1}=e_{1}$, we have that $d^{\prime}_{0}=d_{0}$, so $f(d^{\prime}_{0})=f(d_{0})=f(d)$.
Proof A.13 (Proof of Lemma 4.4).
The algorithm performs a depth-first search on the local collapsible edges. It removes the edges of the depth-first search forest and makes each edge adjacent to a vertex in the forest now be adjacent to the root of a tree in the forest, using the map $\varphi$. This is equivalent to collapsing the forest, which produces a homology-equivalent graph, by Lemma 4.3. To see that the output contains no local collapsible edges, note that all local collapsible edges in the initial graph are identified by the algorithm, and collapsing an edge cannot make any previously non-collapsible edge be collapsible, by Lemma 3.5(1). Moreover, since collapsing a local collapsible edge does not change the $f$-value of the endpoints of any edge, a local collapse cannot make a previously non-local, collapsible edge become local collapsible.
The time complexity of the depth-first search is linear.
Proof A.14 (Proof of Proposition 4.5).
The guiding principle for collapsing collapsible edges followed by the algorithm is simple: first identify a maximal monotonic tree (Definition 4.2) of collapsible edges; then collapse that subtree, which does not change the homology by Lemma 4.3; and then iterate this. In order to do this efficiently, note that any maximal monotonic tree of collapsible edges contains exactly one local connected component of the graph (that is, a connected component in the subgraph of vertices and edges having the same $f$-value) that is minimal (that is, that has no adjacent collapsible edge decreasing the grade). Since the input is assumed to not have any locally collapsible edges, minimal local components are the same as minimal vertices, which are easy to identify, since these are simply vertices with no adjacent collapsible edge decreasing the grade.
The algorithm iterates over minimal vertices, and collapses the corresponding DFS tree of collapsible edges going up in the filtration. Since collapsing edges cannot produce more collapsible edges (Lemma 3.5(1)), nor more minimal vertices, at the end of the main for-loop there are no collapsible edges left.
The output Betti table $\beta_{0}$ is correct by Theorem 3.8(1), and the time complexity of the depth-first search is linear.
Proof A.15 (Proof of Proposition 4.6).
We start with the claim about the time complexity. Algorithm 2 is linear by Proposition 4.5, so it remains to prove that we can check the condition $[e_{0}]=[e_{1}]\in\pi_{0}(V_{e},E_{e},\partial^{\prime})$ in $O(|G|)$ time. To do this, we run breadth-first search on $G^{\prime}$, starting from $e_{0}$, and simply ignore every vertex and every edge whose $f$-value is not less than or equal to $f(e)$; and we also ignore $e$. If at some point we encounter $e_{1}$, then $[e_{0}]=[e_{1}]\in\pi_{0}(V_{e},E_{e},\partial^{\prime})$, and otherwise $[e_{0}]\neq[e_{1}]\in\pi_{0}(V_{e},E_{e},\partial^{\prime})$.
We now prove the correctness of the algorithm. By Theorem 3.8, we only need to check that, at the end of the for-loop in Line 3, the graph $(V^{\prime},E^{\prime},\partial^{\prime},f)$ contains no deletable edges. Clearly, the for-loop visits all edges, so correctness follows from Lemma 3.5(2).
A.4 Proofs of main theorems
See A
Proof A.16.
This follows directly from Propositions 4.5 and 4.6.
See B
Proof A.17.
The output $\beta_{0}$ is correct by the correctness of Algorithm 2. To prove correctness of the rest of the algorithm, we observe the following: by iterating over edges one-by-one in a way that respects the lexicographic order, we are in fact iterating over edges according to a total order $\prec$ on $E^{\prime}$ that refines the lexicographic order induced by $f$. We consider the following invariant:
  • ()($\ast$)( ∗ )
    At the end of the iteration of the main for-loop in Line 4 corresponding to the edge $e\in E^{\prime}$, the dynamic dendrogram $D$ is a dynamic dendrogram for the filtered graph $(V^{\prime},\{e^{\prime}\in E^{\prime}:e^{\prime}\preceq e\},\partial^{\prime},g)$, where $g(v)\coloneqq-\infty$ for all $v\in V^{\prime}$, and $g(e^{\prime})\coloneqq f_{\mathbf{y}}(e^{\prime})$.
It is clear that the algorithm maintains invariant ()($\ast$)( ∗ ), since $\prec$ is, by definition, the total order in which we visit the edges.
We now address the correctness of the outputs $\beta_{0}^{1}$, $\beta_{1}$, and $\mathcal{M}$. An element of $\beta_{0}^{1}$ is added either in the if-clause in Line 4 checking for $s\leq y$, or when an element of $\beta_{2}$ is added in Line 4. The if-clause corresponds exactly to an edge that is connecting two vertices in the same connected component, that is, to a deletable edge. Thus, by Theorem 3.10 and Lemma 3.7, the output $\beta_{0}^{1}$ is correct as long as the output for $\beta_{2}$ is. It also follows that the output $\beta_{1}$ is correct, as deletable edges are being identified correctly, and that the output presentation matrix $\mathcal{M}$ is correct, by Theorem 3.8.
To prove that the output $\beta_{2}$ is correct, we need to show that cycle-creating edges are being identified correctly. To see this, note that the check $s<\infty$ is being performed only after knowing that the edge $e$ is not deletable (and thus after knowing that the edge will be part of the minimal graph, by Lemma 3.5), and that the if-clause $s<\infty$ is true exactly when $e_{0}$ and $e_{1}$ belong to the same connected component at some grade $(x^{\prime},y^{\prime})$ with $x^{\prime}<f_{\mathbf{x}}(e)$, by the invariant ()($\ast$)( ∗ ). This is precisely the case when $e$ is cycle-creating, by Definition 3.9, and we have $s=f_{\mathbf{y}}(w_{e})$, where $w_{e}$ is a minimal witness of $e$, also by the invariant ()($\ast$)( ∗ ). If follows that the output $\beta_{2}$ is correct.
To conclude, we analyze the time complexity. Since the algorithm consists of for-loops visiting each vertex and each edge exactly once, we only need to make sure that the operations involving $D$ are in $\log(|G|)$ which is proven in [21, Section 4].
Appendix B Multi-critical filtrations
If $X$ is a set, let $\mathsf{Parts}_{f}(X)$ denote the set of finite, non-empty subsets of $X$. A finite multi-critical filtration of a graph $G=(V,E)$ by a poset $\mathscr{P}$ consists of functions $f_{V}:V\xrightarrow{\;\;\;}\mathsf{Parts}_{f}(\mathscr{P})$ and $f_{E}:E\xrightarrow{\;\;\;}\mathsf{Parts}_{f}(\mathscr{P})$ such that, for every $e\in E$ and $p\in f_{E}(e)$ there exists $q_{0}\in f_{V}(e_{0})$ and $q_{1}\in f_{V}(e_{1})$ with $q_{0},q_{1}\leq p$.
Let $\mathscr{P}$ be a lattice, and let $(G,f)$ be a finite multi-critical filtration. There exist algorithms for reducing the computation of the homology of multi-critical filtrations to that of critical filtrations [14], but these cannot, a priori, be directly combined with our algorithms. We thus describe a custom procedure for constructing a ($1$-critical) filtered graph with the same 00-dimensional persistent homology as $(G,f)$.
In words, we proceed as follows: for each vertex in the original filtration, we add a new copy of the vertex at all the grades at which the vertex is born and an edge between copies of the same vertex at all the minimal joins of the birth grades; and for each edge in the original filtration, we add a new copy of the edge (connecting any one copy of its endpoints) at all the grades at which the edge is born. More precisely, we proceed as follows. As set of vertices, we take
$V^{\prime}\coloneqq\{(v,p):v\in V,p\in f_{V}(v)\}.$
In order to identify all the different copies of a vertex $v$, we first define
$E^{\prime\prime}_{v}\coloneqq\{(v,p,p^{\prime}):p\neq p^{\prime}\in f_{V}(v)\}\;\text{ with }\;\partial^{\prime}(v,p,p^{\prime})=((v,p),(v,p^{\prime})).$
Consider the function $\varphi:E^{\prime\prime}_{v}\xrightarrow{\;\;\;}\mathscr{P}$ given by $\varphi(v,p,p^{\prime})=p\vee p^{\prime}$. We then let $E^{\prime}_{v}\subseteq E^{\prime\prime}_{v}$ be such that the cardinality of $E^{\prime}_{v}$ is minimal and $\varphi(E^{\prime}_{v})$ contains all the minimal elements of $\varphi(E^{\prime\prime}_{v})$. The idea is that, in order to identify the different copies of a vertex $v$, it is sufficient to use the edges in $E^{\prime}_{v}$. To account for the edges of $G$, let
$E^{\prime}_{edg}\coloneqq\{(e,p):e\in E,p\in f_{E}(e)\}\;\text{ and }\;\partial^{\prime}(e,p)=((e_{0},q_{0}),(e_{1},q_{1})),$
where $q_{0}\in f_{V}(e_{0})$ and $q_{1}\in f_{V}(e_{1})$ satisfy $q_{0},q_{1}\leq p$. As set of edges, we take $E^{\prime}\coloneqq\left(\bigcup_{v\in V}E^{\prime}_{v}\right)\cup E^{\prime}_{edg}$, and define the filtration as follows
$f_{V^{\prime}}(v,p)=p,\;\;f_{E^{\prime}}(v,p,p^{\prime})\coloneqq p\vee p^{\prime}\,\,,\,\text{ and }\;f_{E^{\prime}}(e,p)\coloneqq p.$
We leave it to the reader to check that the 00-dimensional persistent homology of the $1$-critical filtered graph $(V^{\prime},E^{\prime},\partial^{\prime},f_{V^{\prime}},f_{E^{\prime}})$ is isomorphic to that of $(G,f)$. One way to see this, is to show that $\pi_{0}(V^{\prime},E^{\prime},\partial^{\prime},f_{V^{\prime}},f_{E^{\prime}})\cong\pi_{0}(G,f)$ using the continuous function from the geometric realization of $(V^{\prime},E^{\prime},\partial^{\prime})$ to that of $(V,E)$ defined by $|(v,p)|\mapsto|v|$, $|(v,p,p^{\prime})|\mapsto|v|$, $|(e,p)|\mapsto|e|$.